OpenStudy (anonymous):

x 2 - y = 3 x - y = -3

OpenStudy (anonymous):

-y=-x-3 y=x+3 x^2-(x+3)=3 x^2-x+3=3 x^2-x=0 idk the rest of it :/ sorry